An area of a rectangle is 108. The ratio of the Width to Length is 3:4. What is the Length and the Width.
Answer by RAY100(1637)

Area = Length * Width = 108
.
w/L = 3/4,,,,4w=3L,,,,w=3L/4,,,,subst
.
L*(3L/4)=108
.
(3/4)*L^2 =108
.
L^2 = 144
.
L=12,,,,,,W=3/4 *12 = 9
.
check,,,,12*9=108,,,,ok